The mass (M) and the radius (r) of a nucleus can be expressed in terms of the mass number, A. (a) Show that the density of a nucleus is independent of A. (b) Calculate the density of a gold (Au) nucleus. Compare your answer to that for iron (Fe). 


1
Expert's answer
2021-06-29T02:06:02-0400

Mass=M

Radius=r

Mass number =A

Part(a)

"\rho=\frac{M}{V}=\frac{M}{\frac{4 \pi r^3}{3}}"

"r=r_0A^\frac{1}{3}"

"\rho=\frac{M}{4\pi r_0^3 A}"

Part (b)


"\rho=\frac{1.67\times10^{-27}}{4\times \pi 4.2\times10^{-15}\times{197}}=19.3g/cm^3"

Iron

"\rho=7.87gm/cm^3"

Difference density of iron and gold

"\rho=19.3-7.87=11.7g/cm^3"